OpenStudy (anonymous):

The perimenter of a rectangle is 59 feet.If the legth is 14 feet,find the width

OpenStudy (anonymous):

14+14=28 28-59=31 31/2=15.5

OpenStudy (anonymous):

width is 15.5 feet on each side